OpenStudy (anonymous):

Simplify the following expression : √16/2+6 to the second power √38 38 40 100

OpenStudy (anonymous):

\[\frac{\sqrt{16}}{2}+6^2=\sqrt{\frac{16}{4}}+36=\sqrt{4}+36=38\]